I need help to solve this problem and graph it. y=(x-2)^2+6
Answer by mananth(16949)

y=(x-2)^2+6
y= x^2-4x+4+6
y=x^2-4x+10
x=0, y=10 (0,10)
x=2,y=6 (2,6)
x=3,y=7 (3,7)
x=-1, y=15 (-1,15)
x=-2,y=22 (-2,22)
Since this is a quadratic equation it is a parabola.
since co-efficient of x^2 is positive Then the graph opens upwards